OpenStudy (anonymous):

Write an equation for the line that is parallel to the given line and that passes through the given point. y = 5/4x – 10; (–6, –29) @Hero

OpenStudy (anonymous):

i have a way i did it but i wanna know if its right

hero (hero):

What have you done so far?

OpenStudy (anonymous):

i just plugged the given points where the ten is but then i got stuck

hero (hero):

Can you show what you have done more explicitly?

OpenStudy (anonymous):

y-5/4x+-6 for the first point

hero (hero):

Are you sure you know what you're doing?

OpenStudy (anonymous):

the same but with -29 for the second

OpenStudy (anonymous):

and no i dont lol thats why im asking for help ;/

hero (hero):

1st Observation: The general form of the equation of a line is y = mx + b 2nd Observation: The equation of the given line is y = (5/4)x - 10 3rd Observation Both lines are parallel to each other which means their slopes are the same; therefore, the equation of the line parallel to the given line (so far) is: y = (5/4)x + b 4th Observation The point (x,y) = (-6,-29) is given, therefore substitute the point into the equation of the parallel line to find b: -29 = (5/4)(-6) + b Finish solving for b.

OpenStudy (anonymous):

b=−43/2

hero (hero):

I get b = -23/2 Double check your work.

hero (hero):

When you do, let me know what the equation of the parallel line is.

OpenStudy (anonymous):

okay i got it